Operation Summary for July 4, 2011 - July 6, 2011 |
Monday July 4 The day shift began with Operations monitoring the Tevatron (TEV) store 8858 (98.02E30) and the Recycler stash (213.4E10), with the Antiproton Source (Pbar) stacking (23.5E10), with the Main Injector (MI), Linac, and Booster operating normally, with the NuMI, MiniBooNE, and Meson FTBF T-1015 experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams. During the day shift, Operations terminated store 8858 at 12:07 PM. Operators prepared D0 and CDF for access at 12:22 PM; both experiments were out of access by 2:01 PM. Operators put the TEV through a wet squeeze at 3:28 PM. Operations put the TEV into shot setup at 3:48 PM. Operators reported at 3:52 PM that one of the air conditioning units in the MI-12 service building tripped off; the Duty electrician investigated. Operators reported a short downtime due to a Booster RF station (BRF4). During the evening shift, Operations established store 8861 at 4:56 PM with an average initial luminosity (AIL) of 381.75E30. Operations disabled beam to T-1015 at 6:39 PM. Tuesday July 5 The midnight shift began with Operations monitoring store 8861 (72.83E30) and the stash (378.8E10), with Pbar stacking (5.4E10), and with NuMI and MiniBooNE taking beam. During the midnight shift, Pbar lost its beam permit at 3:55 AM due to beam position problems; operators investigated and loaded an old file into Pbar’s autotune program. Operations established beam to T-1015 at 4:14 AM, but two hours later Operations disabled beam so that experimenters could access to investigate equipment problems. During the day shift, Operations terminated store 8861 at 8:40 AM. Cryo system technicians began work on tightening slipping belts for the TEV sectors A1, A2, and A3 wet engines. Operations resumed sending beam to T-1015 at 10:04 AM and then shut them off about 20 minutes later. Operations put the TEV into shot setup at 10:12 AM and established store 8862 at 12:22 PM with AIL of 344.12E30. MTA began taking beam. Operators began preparing Booster for access at 12:38 PM for RF work. Operators prepared MI, Recycler, and the Pelletron for access at 12:49 PM. Operators prepared MTA for access at 1:45 PM. Operators prepared MI-65 at 1:52 PM. A Water Works person reported at 2 PM that there was no flow of pond water through the B4 pond pump; He and FESS personnel cleaned the filter at B4, but it didn’t help. They thought the heat exchanger was plugged. Also, the D3 LCW pump would not turn on; their investigations continued. Operators reported at 3:59 PM that the D3 LCW temperatures were coming down. Operators reported short downtimes due to a MI power supply trip on a MI-40 lower filter indication, and due to trips on LRF2 and two kickers (KPS5S and KPS4). During the evening shift, Booster came out of access at 4:40 PM. MI came out of access at 5:15 PM. Operations had beam to MiniBooNE, NuMI, and Pbar by 6:10 PM. Operators investigated the high LCW temperatures at C1 & 2 and discovered a closed three-way valve; they put the valve in its proper position and temperatures came down. Operators performed a MI coalescing check at 11:04 PM. At 11:18 PM, operators reported a NuMI evaporator alarm; they contacted a NuMI person. Wednesday July 6 The midnight shift began with Operations monitoring store 8862 (87.92E30) and the stash (170.9E10), with Pbar stacking (23.3E10), and with NuMI and MiniBooNE taking beam. During the midnight shift, at 2:34 AM, a MI kicker (KPS5S) trip caused the loss of antiprotons (8 mA) during a transfer. It tripped off several more times and then would not reset. Operators called in Kicker personnel at 3:51 AM. They brought the kicker back online and Pbar resumed stacking at 5:18 AM. Unfortunately, KPS5S tripped off again at 5:25 AM. Kicker personnel reported that the kicker’s Thyratron tube needed replacing. Operators masked a bit on KPS5S at 5:36 AM so that NuMI could take beam. Pbar resumed stacking at 6:39 AM. The Plans for Wednesday The plans for today are to terminate store 8862 this morning and allow CDF and D0 a two-hour access. MI will be down for six to eight hours while Kicker personnel replace the Thyratron tube for KPS5S. Interlock personnel will also conduct Pbar safety system testing and, if there is time, TEV safety system testing. |
|
More Information Comments and Suggestions |
| last modified 07/06/2011 email Fermilab |
| Security, Privacy, Legal | |